You are working on a research project in which you must control the direction of travel of electrons using deflection plates.Can someone please help me with this? I'll rate you if everything is correct. I'm not sure if my first part is correct, but PLEASE do all parts
Answer #1

Given; Length of the plate is L = 0.25m Separation distance between the plates is d=1.68 cm Electrons are fired at the speedL y d 2 Ē Vi + + + + + + + +(a) The electric field acting on the electron inside he apparatus is F =-eĒ F = -eEųThe acceleration of the electron inside the apparatus ma = -eEy E â= m(t) = Ū; + āt t(t) = 0; Cos 6,4 + ; sin 6 g Et -ý m e Et (t) = 0; cos 0; î + ; sin ; CE ý (1) m The x-xomponent of the velociEquation (1) can be written as follows üſt = = 0; cos 0 t + i sin 0; eEl mu, cos y U; cos,The angle at which electron leaves the apparatus with respect to x-axis is given by tan e El tan v; sin ; mu; cose; U; cose,Minimum angle occurs when the electron leaves just above the right corner of the lower plate. Net displacement along the y-axSubstitute E into the equation(1),we have X tan min tanmin el Vi sin , 2mo sin 0; cos; mo; cos, el U; cose; 0; sin 0; -20; si0; = tan (6) =tan-1 1.68 x 10-2 0.250 0; = 3.43° Hence, the range of angle is from min = -3.84 to Omar = 3.84°(b) The electric field require for maximum possible deviation is 2mv sin 6; cose E= el E mv sin 20; el 9.11 x 10-31 x (5.1x 1(a) The range of angle is from min = -3.84º to Omar = 3.84 (b) The electric field require for maximum possible deviation is 7

I hope you understood the problem and got your answers, If yes rate me!! or else comment for a better solutions.